the length of a rectangular field is 7 meters more than twice it width. the perimeter is 140 meters, find the dimension of the field
Answer by checkley77(12844)

L=2W+7
2L+2W=140
2(2W+7)+2W=140
4W+14+2W=140
6W=140-14
6W=126
W=126/6
W=21 ANS. FOR THE WIDTH.
L=2*21+7=42+7=49 ANS. FOR THE LENGTH.
PROOF:
2*49+2*21=140
98+42=140
140=140